In a location where the speed of sound in air is 354ms-1, a 200Hz sound wave impinges on two slits 30.0cm apart.

           a.) At what angle is the first maximum located?

           b.) If the sound is replaced by a 3.00cm microwave, what slit separation gives the same angle for the first maximum?


1
Expert's answer
2021-02-22T10:27:19-0500

Answer

(A)


"\lambda=\frac{v}{f}"

"\lambda=\frac{354}{200}=1.77m"


Angle can be obtain as

"\theta=sin^{-1}(\frac{m\lambda}{d})"


By putting value


"\theta=sin^{-1}(\frac{(1)(0.177)}{0.30})=36.1^°"


(B)

"d=\frac {m\lambda}{sin\theta}"

By putting value


"d=\frac {(1)(0.0300)}{sin(36.1)°}"


"d=0.0508m"